# A HHO formulation for variable density incompressible flows where the density is purely advected

### v1

@@ Abstract
We propose a Hybrid High-Order (HHO) formulation of the incompressible Navier–Stokes equations with variable density that provides exact conservation of volume and, accordingly, pure advection of the density variable. The spatial discretization relies on hybrid velocity-density-pressure spaces and the temporal discretization is based on Explicit Singly Diagonal Implicit Runge-Kutta (ESDIRK) methods. The formulation possesses some attractive features that can be fruitfully exploited for the simulation of mixtures of immiscible incompressible fluids, namely: pressure-robustness, conservation of volume enforced cell-by-cell up to machine precision, robustness in the convection dominated regime thanks to upwind stabilizations, weak imposition of boundary conditions, implicit high-order accurate time stepping encompassing unsteady boundary conditions, reduced memory footprint thanks to static condensation, possibility to exploit inherited $p$-multilevel solution strategies to improve performance of iterative solvers. Numerical validation is performed showcasing convergence rates, pressure-robustness and pure advection of the density, thus we tackle the Rayleigh-Taylor instability at Reynolds 10001~000 and 50005~000.
@@ 1 Introduction
In this work we extend the HHO formulations of [7]-[8], devoted to constant density incompressible flows, to cope with mixtures of immiscible incompressible fluids with variable density. The present research effort relies on a hybrid discretization of all the variables of the problem, including velocity, pressure and density. The choice of a hybrid pressure, combined with an appropriate selection of local polynomial spaces, allows to reconstruct an $\bm{H}(\operatorname{div})$-conforming velocity approximation starting from non-conforming polynomial spaces. Thanks to $\bm{H}(\operatorname{div})$-conformity and exact incompressibility, the mass conservation equation can be formulated as the material time derivative of the density, both at the continuous and at the discrete level. Moreover, the HHO discretization of mass transport uses to good advantage the availability of a solenoidal velocity field, see e.g., Di Pietro et al. [13]. Pressure-robustness implies that irrotational body forces only affect the pressure field and, as a consequence, the velocity and density errors are insensitive to the accuracy of the pressure field representation. Robustness in the convection-dominated regime is achieved by introducing upwinding stabilizations for the transport of mass and momentum, in tandem with suitable discretization of the density time derivative.
@@
In order to treat time-dependent problems, we rely on high-order accurate Explicit Singly Diagonal Implicit Runge-Kutta (ESDIRK) time integration schemes especially conceived for Differential Algebraic Equations (DAEs) of differential index equal to two. Efficient resolution of the algebraic problems resulting from the space-time discretization relies on $p$-multigrid preconditioners in the spirit of Botti et al. [9], specifically conceived to work in tandem with Schur complement solution strategies. Thanks to static condensations, the blocked structure, the dimension of the blocks and, ultimately, the sparsity of the global matrix resulting from HHO discretizations is dictated by face unknowns. Interestingly, since the dimension of d1d{-}1-variate polynomial spaces grows slower that the dimension of $d$-variate polynomial spaces when increasing $k$, both the memory footprint of the global matrix and the cost of matrix-vector products involved in iterative solution strategies benefit from static condensation.
@@
Several methods for variable density incompressible flows have been proposes in literature. A DG formulation based on exact Riemann solvers for the artificial compressibility perturbation of the problem was proposed by Bassi et al. [3] and assessed in [2]. An entropy–stable DG approximation was devised by Manzanero et al. [21] while a post-processed semi-implicit DG scheme was introduced by [20]. A Discrete Duality Finite Volume (DDFV) scheme was proposed by Goudon et al. [15]. Recently, Fu [14] addressed the Cahn-Hilliard phase-field model with a HDG formulation based on divergence-free velocity spaces. Earlier works focused on projection methods, consider for example Guermond et al. [16]-[17].
@@
The paper is organized as follows. In Sections 2 and 3 we respectively introduce the model problem and present its discrete formulation. $\bm{H}(\operatorname{div})$-conformity and pressure robustness are addressed in Section 3.5. Numerical validation of spatial and temporal convergence rates is performed in Section 4.1 and Section 4.2 considering exact and manufactured solutions. A realistic flow configuration is tackled in Section 4.3 where we consider a low Atwood number Rayleigh-Taylor instability at $\mathrm{Re}{=}1~000$ and $\mathrm{Re}{=}5~000$. Finally, some conclusions are drawn in Section 5.
@@ 2 Continuous setting
Equations (1a), (1b) and (1c) are the linear momentum, the mass conservation and the volume conservation equations, respectively. We remark that the discrete formulation will be designed mimicking as closely as possible the weak formulation of the continuous problem. Accordingly, conservation of volume will be exactly enforced at the discrete level, ensuring that the density is a purely advected quantity.
@@ 3.1 Mesh
The HHO formulation proposed in this work is capable of dealing with two and three dimensional flow problems. Nevertheless, since numerical test cases focus on two space dimensions, discrete settings are provided for the 2D case focusing on simplicial meshes. We consider meshes of the domain $\Omega$ corresponding to couples $\mathcal{M}_{h}\coloneq(\mathcal{T}_{h},\mathcal{F}_{h})$, where $\mathcal{T}_{h}$ is a finite collection of triangular elements such that $h\coloneq\max_{T\in\mathcal{T}_{h}}h_{T}>0$ with $h_{T}$ denoting the diameter of $T$, while $\mathcal{F}_{h}$ is a finite collection of line segments, with $h_{F}$ denoting the length of $F$. It is assumed henceforth that the mesh $\mathcal{M}_{h}$ is shape- and contact-regular, as detailed in [12, Definition 1.4] and that its trace on $\partial\Omega$ is compatible with the partition $\partial\Omega=\partial\Omega_{\rm D}\sqcup\partial\Omega_{\rm N}$. We respectively denote by $\mathcal{F}_{h}^{\rm D}$ and $\mathcal{F}_{h}^{\rm N}$ the sets of Dirichlet and Neumann faces, whereas the set of internal faces is $\mathcal{F}_{h}^{\rm i}\coloneq\mathcal{F}_{h}\setminus\big(\mathcal{F}_{h}^{\rm D}\cup\mathcal{F}_{h}^{\rm N}\big)$. For each mesh cell $T\in\mathcal{T}_{h}$, the faces contained in the cell boundary $\partial T$ are collected in the set $\mathcal{F}_{T}$ and we additionally let $\partial T^{\bullet}\coloneqq\partial T\cap\partial\Omega_{\bullet}$ for $\bullet\in\{{\rm D},{\rm N}\}$. For all $T\in\mathcal{T}_{h}$ and all $F\in\mathcal{F}_{T}$, $\bm{n}_{TF}$ denotes the unit vector normal to $F$ pointing out of $T$ and $\bm{n}_{\partial T}:\partial T\to\mathbb{R}^{d}$ is such that $\bm{n}_{\partial T}|_{F}\coloneqq\bm{n}_{TF}$ for all $F\in\mathcal{F}_{T}$.
@@ 3.2 Discrete spaces
$\underline{Z}_{h}^{k}\coloneqq\left\{\underline{z}_{h}=\big((z_{T})_{T\in\mathcal{T}_{h}},(z_{F})_{F\in\mathcal{F}_{h}}\big)\,:\,\begin{array}[]{l}\text{$z_{T}\in\mathcal{P}^{k}(T)$ for all $T\in\mathcal{T}_{h}$},\\ \text{$z_{F}\in\mathcal{P}^{k}(F)$ for all $F\in\mathcal{F}_{h}$}\end{array}\right\},$
$\underline{Q}_{h}^{k+1}\coloneqq\left\{\underline{q}_{h}=\big((q_{T})_{T\in\mathcal{T}_{h}},(q_{F})_{F\in\mathcal{F}_{h}}\big)\,:\,\begin{array}[]{l}\text{$q_{T}\in\mathcal{P}^{k}(T)$ for all $T\in\mathcal{T}_{h}$},\\ \text{$q_{F}\in\mathcal{P}^{k+1}(F)$ for all $F\in\mathcal{F}_{h}$}\end{array}\right\}.$
@@ 3.4 Semi-discrete HHO formulation
Given $(\underline{\bm{u}}_{T},\underline{\rho}_{T},\underline{p}_{T})\in\underline{\bm{V}}_{T}^{k}\times\underline{Z}_{T}^{k}\times\underline{Q}_{T}^{k+1}$, the (local) spatial discretizations $sd^{\rm slm}_{T,{\rm!bc}}((\underline{\bm{u}}_{T},\underline{\rho}_{T},\underline{p}_{T});\cdot):\underline{\bm{V}}_{T}^{k}\to\mathbb{R}$ of the steady momentum equation, $sd^{\rm sms}_{T,{\rm!bc}}(\underline{\bm{u}}_{T},\underline{\rho}_{T};\cdot):\underline{Z}^{k}_{T}\to\mathbb{R}$ of the steady mass conservation equation and $sd^{\rm vol}_{T,{\rm!bc}}(\underline{\bm{u}}_{T};\cdot):\underline{Q}^{k+1}_{T}\to\mathbb{R}$ of the volume conservation equation are such that, for all $\underline{\bm{v}}_{T}\in\underline{\bm{V}}_{T}^{k}$, all $\underline{z}_{T}\in\underline{Z}_{T}^{k}$ and all $\underline{q}_{T}\in\underline{Q}_{T}^{k+1}$,
$sd^{\rm slm}_{T,{\rm!bc}}$ $\bigl((\underline{\bm{u}}_{T},\underline{\rho}_{T},\underline{p}_{T});\underline{\bm{v}}_{T}\bigr)=-\int_{T}\rho_{T}(\bm{u}_{T}\otimes\bm{u}_{T}):\nabla\bm{\pi}_{T}^{k}\bm{v}_{T}$ (2)
$+\int_{\partial T}\left[\bigl(\bm{u}_{T}{\cdot}\bm{n}_{\partial T}\bigr)^{\oplus}{\rho}_{T}\bm{u}_{T}{+}\bigl(\bm{u}_{T}{\cdot}\bm{n}_{\partial T}\bigr)^{\ominus}\rho_{\partial T}\bm{u}_{\partial T}\right]\cdot(\bm{\pi}_{T}^{k}\bm{v}_{T}-\bm{\pi}_{\partial T}^{k}\bm{v}_{\partial T})$ (3)
$+\int_{T}\mu\,{{\mathcal{G}}_{T}^{k}}\underline{\bm{u}}_{T}:{{\mathcal{G}}_{T}^{k}}\underline{\bm{v}}_{T}+\int_{\partial T}\frac{\mu}{h_{T}}\,\bm{\pi}_{\partial T}^{k}(\bm{u}_{T}-\bm{u}_{\partial T})\cdot\bm{\pi}_{\partial T}^{k}(\bm{v}_{T}-\bm{v}_{\partial T})$ (4)
@@
$sd^{\rm sms}_{T,{\rm!bc}}$ $\bigl((\underline{\bm{u}}_{T},\underline{\rho}_{T});\underline{z}_{T}\bigr)=$
$-\int_{T}\rho_{T}\bm{u}_{T}\cdot\nabla z_{T}\,+\int_{\partial T}\left[\bigl(\bm{u}_{T}{\cdot}\bm{n}_{TF}\bigr)^{\oplus}{\rho}_{T}{+}\bigl(\bm{u}_{T}{\cdot}\bm{n}_{TF}\bigr)^{\ominus}\rho_{\partial T}\right]({z}_{T}-{z}_{\partial T}),$ (6)
@@
where $\alpha^{\oplus}\coloneqq\frac{1}{2}(\alpha+|\alpha|)$ and $\alpha^{\ominus}\coloneqq\frac{1}{2}(\alpha-|\alpha|)$, $\bm{\pi}_{\partial T}^{k}$ is the $L^{2}$-orthogonal projector onto the broken polynomial space $\mathcal{P}^{k}(\partial T)^{d}$ defined as follows
$\mathcal{P}^{k}(\partial T)^{d}\coloneqq\left\{\bm{v}\in L^{2}(\partial T)^{d}\;:\;\text{$\bm{v}|_{F}\in\mathcal{P}^{k}(F)^{d}$ for all $F\in\mathcal{F}_{T}$}\right\},$
@@
and $\bm{\pi}_{T}^{k}$ is the $L^{2}$-orthogonal projector onto $\mathcal{P}^{k}(T)^{d}$. We remark that upwinding is introduced in the mass conservation and linear momentum equations for enhancing the stability of the discrete formulation. The idea of utilizing $L^{2}$-projections of test functions in the convective term formulation of (2)-(3) is inspired by [4] and improves the accuracy of the method in the convection-dominated regime. The diffusion term stabilization of (4), also involving $L^{2}$-projection operators, was first proposed by [19], see also [11]. Note that the HHO discretization of the mass conservation equation in (6) only relies on the cell velocity $\bm{u}_{T}$ which, as result of the HHO discretization of the volume conservation in (7), is $\bm{H}(\operatorname{div})$-conforming and divergence-free, see Section 3.5 for details. Accordingly, the trace of the normal velocity component is continuous on $\partial T$.
@@
Let’s now take into account Dirichlet and Neumann boundary conditions. Given $(\underline{\bm{u}}_{T},\underline{\rho}_{T},\underline{p}_{T})\in\underline{\bm{V}}_{T}^{k}\times\underline{Z}_{T}^{k}\times\underline{Q}_{T}^{k+1}$, the (local) spatial discretizations $sd^{\rm slm}_{T}((\underline{\bm{u}}_{T},\underline{\rho}_{T},\underline{p}_{T});\cdot):\underline{\bm{V}}_{T}^{k}\to\mathbb{R}$ of the steady momentum equation, $sd^{\rm sms}_{T}(\underline{\bm{u}}_{T},\underline{\rho}_{T};\cdot):\underline{Z}^{k}_{T}\to\mathbb{R}$ of the steady mass conservation equation and $sd^{\rm vol}_{T}(\underline{\bm{u}}_{T};\cdot):\underline{Q}^{k+1}_{T}\to\mathbb{R}$ of the volume conservation equation are such that, for all $\underline{\bm{v}}_{T}\in\underline{\bm{V}}_{T}^{k}$, all $\underline{z}_{T}\in\underline{Z}_{T}^{k}$ and all $\underline{q}_{T}\in\underline{Q}_{T}^{k+1}$,
$sd^{{\rm slm}}_{T}$ $\bigl((\underline{\bm{u}}_{T},\underline{\rho}_{T},\underline{p}_{T});\underline{\bm{v}}_{T}\bigr)\coloneqq sd^{{\rm slm}}_{T,{\rm!bc}}((\underline{\bm{u}}_{T},\underline{p}_{T});\underline{\bm{v}}_{T})$
@@
${+}\int_{\partial T^{\rm N}}\bigl(\bm{u}_{\partial T}{\cdot}\bm{n}_{\partial T}\bigr)^{\oplus}\bigl(\rho_{\partial T}\bm{u}_{\partial T}{\cdot}{\bm{\pi}_{\partial T}^{k}}\bm{v}_{\partial T}\bigr){+}\int_{\partial T^{\rm N}}\bigl(\bm{u}_{\partial T}{\cdot}\bm{n}_{\partial T}\bigr)^{\ominus}\bigl(g_{\rm In}\bm{u}_{\partial T}{\cdot}{\bm{\pi}_{\partial T}^{k}}\bm{v}_{\partial T}\bigr)$
$-\int_{\partial T^{\rm N}}p_{\partial T}(\bm{v}_{\partial T}\cdot\bm{n}_{\partial T})+\int_{\partial T^{\rm N}}\bm{g}_{\rm N}\cdot\bm{v}_{\partial T}$
@@
$+\int_{\partial T^{\rm D}}\bigl(\bm{g}_{\rm D}\cdot\bm{n}_{\partial T}\bigr)^{\oplus}\;(\rho_{\partial T}\bm{u}_{\partial T}\cdot\bm{v}_{\partial T})+\int_{\partial T^{\rm D}}\bigl(\bm{g}_{\rm D}\cdot\bm{n}_{\partial T}\bigr)^{\ominus}\;(g_{\rm In}\bm{g}_{\rm D}\cdot\bm{v}_{\partial T})$

### v3

@@ Abstract
We propose a Hybrid High-Order (HHO) formulation of the incompressible Navier–Stokes equations with variable density that provides exact conservation of volume and, accordingly, pure advection of the density variable. The spatial discretization relies on hybrid velocity-density-pressure spaces and the temporal discretization is based on Explicit Singly Diagonal Implicit Runge-Kutta (ESDIRK) methods. The formulation possesses some attractive features that can be fruitfully exploited for the simulation of mixtures of immiscible incompressible fluids, namely: conservation of volume enforced cell-by-cell up to machine precision, pressure-robustness, ability to preserve density bounds at low-order, robustness in the convection dominated regime, weak imposition of boundary conditions, implicit high-order accurate time stepping, reduced memory footprint thanks to static condensation, possibility to exploit inherited $p$-multilevel solution strategies to improve the performance of iterative solvers. After addressing stability at the discrete level, numerical validation is performed showcasing spatial and temporal convergence rates. To conclude, we tackle the Rayleigh-Taylor instability at different Atwood and Reynolds numbers focusing on mesh independence capabilities.
@@ 1 Introduction
In this work we extend the HHO formulations of [8]-[9], devoted to constant density incompressible flows, to cope with mixtures of immiscible incompressible fluids with variable density. The present research effort relies on a hybrid discretization of all the variables of the problem, including velocity, pressure and density. The choice of a hybrid pressure, combined with an appropriate selection of local polynomial spaces, allows to reconstruct an $\bm{H}(\operatorname{div})$-conforming velocity approximation starting from non-conforming polynomial spaces. Thanks to $\bm{H}(\operatorname{div})$-conformity and exact incompressibility, the mass conservation equation can be formulated as the material time derivative of the density, both at the continuous and at the discrete level. Moreover, the HHO discretization of mass transport uses to good advantage the availability of a solenoidal velocity field, see e.g., Di Pietro et al. [14]. Pressure-robustness implies that irrotational body forces only affect the pressure field and, as a consequence, the velocity and density errors are insensitive to the accuracy of the pressure field representation. Robustness in the convection-dominated regime is achieved by introducing upwinding stabilizations for the transport of mass and momentum, and suitable time derivative stabilizations for the density variable.
@@
In order to treat time-dependent problems, we rely on high-order accurate Explicit Singly Diagonal Implicit Runge-Kutta (ESDIRK) time integration schemes especially conceived for Differential Algebraic Equations (DAEs) of differential index equal to two. Efficient resolution of the algebraic problems resulting from the space-time discretization relies on $p$-multigrid preconditioners in the spirit of Botti et al. [7], designed to work in tandem with Schur complement solution strategies. Thanks to static condensations, the blocked structure, the dimension of the blocks and, ultimately, the sparsity of the global matrix resulting from HHO discretizations is dictated by face unknowns. Interestingly, since the dimension of d1d{-}1-variate polynomial spaces grows slower that the dimension of $d$-variate polynomial spaces when increasing $k$, both the memory footprint of the global matrix and the cost of matrix-vector products involved in iterative solution strategies benefit from static condensation.
@@
Several methods for variable density incompressible flows have been proposes in literature. A DG formulation based on exact Riemann solvers for the artificial compressibility perturbation of the problem was proposed by Bassi et al. [3] and assessed in [2]. An entropy–stable DG approximation was devised by Manzanero et al. [22] while a post-processed semi-implicit DG scheme was introduced by [21]. A Discrete Duality Finite Volume (DDFV) scheme was proposed by Goudon et al. [16], while Fu [15] addressed the Cahn-Hilliard phase-field model with a HDG formulation based on divergence-free velocity spaces. Recently, Nordström et al. [23] proposed an energy stable formulation for multi-phase incompressible flows and Dauphin et al. [12] analysed a low-order hybrid method for the variable-density incompressible Navier-Stokes equations. Earlier pioneering works focused on Galerkin discretizations with projection methods, consider for example Guermond et al. [17]-[18].
@@
The paper is organized as follows. In Sections 2 and 3 we respectively introduce the model problem and present its discrete formulation. $\bm{H}(\operatorname{div})$-conformity and pressure robustness are addressed in Section 3.5. Numerical validation of spatial and temporal convergence rates is performed in Section 4.1 and Section 4.2 considering exact and manufactured solutions. A realistic flow configuration is tackled in Section 4.3 where we consider a low Atwood number Rayleigh-Taylor instability at $\mathrm{Re}{=}1~000$ and $\mathrm{Re}{=}5~000$ and a high Atwood number Rayleigh-Taylor instability at $\mathrm{Re}{=}1~000$. Finally, some conclusions are drawn in Section 5.
@@ 2 Continuous setting
Equations (1a), (1b) and (1c) are the linear momentum, the mass conservation and the volume conservation equations, respectively. We remark that conservation of volume will be exactly enforced at the discrete level, ensuring that the density is a purely advected quantity.
@@ 3.1 Mesh
The HHO formulation proposed in this work is capable of dealing with two and three dimensional flow problems. Nevertheless, since numerical test cases focus on two space dimensions, discrete settings are provided for the 2D case focusing on simplicial meshes. We consider meshes of the domain $\Omega$ corresponding to couples $\mathcal{M}_{h}\coloneq(\mathcal{T}_{h},\mathcal{F}_{h})$, where $\mathcal{T}_{h}$ is a finite collection of triangular elements such that $h\coloneq\max_{T\in\mathcal{T}_{h}}h_{T}>0$ with $h_{T}$ denoting the diameter of $T$, while $\mathcal{F}_{h}$ is a finite collection of line segments, with $h_{F}$ denoting the length of $F$. It is assumed henceforth that the mesh $\mathcal{M}_{h}$ is shape- and contact-regular, as detailed in [13, Definition 1.4] and that its trace on $\partial\Omega$ is compatible with the partition $\partial\Omega=\partial\Omega_{\rm D}\sqcup\partial\Omega_{\rm N}$. We respectively denote by $\mathcal{F}_{h}^{\rm D}$ and $\mathcal{F}_{h}^{\rm N}$ the sets of Dirichlet and Neumann faces, whereas the set of internal faces is $\mathcal{F}_{h}^{\rm i}\coloneq\mathcal{F}_{h}\setminus\big(\mathcal{F}_{h}^{\rm D}\cup\mathcal{F}_{h}^{\rm N}\big)$. For each mesh cell $T\in\mathcal{T}_{h}$, the faces contained in the cell boundary $\partial T$ are collected in the set $\mathcal{F}_{T}$ and we additionally let $\partial T^{\bullet}\coloneqq\partial T\cap\partial\Omega_{\bullet}$ for $\bullet\in\{{\rm D},{\rm N}\}$. For all $T\in\mathcal{T}_{h}$ and all $F\in\mathcal{F}_{T}$, $\bm{n}_{TF}$ denotes the unit vector normal to $F$ pointing out of $T$ and $\bm{n}_{\partial T}:\partial T\to\mathbb{R}^{d}$ is such that $\bm{n}_{\partial T}|_{F}\coloneqq\bm{n}_{TF}$ for all $F\in\mathcal{F}_{T}$.
@@ 3.2 Discrete spaces
$\underline{Z}_{h}^{k}\coloneqq\left\{\underline{z}_{h}=\big((z_{T})_{T\in\mathcal{T}_{h}},(z_{F})_{F\in\mathcal{F}_{h}}\big)\,:\,\text{$z_{T}\in\mathcal{P}^{k}(T)$ for all $T\in\mathcal{T}_{h}$},\;\text{$z_{F}\in\mathcal{P}^{k}(F)$ for all $F\in\mathcal{F}_{h}$}\right\},$
$\underline{Q}_{h}^{k+1}\coloneqq\left\{\underline{q}_{h}=\big((q_{T})_{T\in\mathcal{T}_{h}},(q_{F})_{F\in\mathcal{F}_{h}}\big)\,:\,\text{$q_{T}\in\mathcal{P}^{k}(T)$ for all $T\in\mathcal{T}_{h}$},\;\text{$q_{F}\in\mathcal{P}^{k+1}(F)$ for all $F\in\mathcal{F}_{h}$}\right\}.$
@@ 3.4 Semi-discrete HHO formulation
Given $(\underline{\bm{u}}_{T},\underline{\rho}_{T},\underline{p}_{T})\in\underline{\bm{V}}_{T}^{k}\times\underline{Z}_{T}^{k}\times\underline{Q}_{T}^{k+1}$, the (local) spatial discretizations $sd^{{\rm slm},\pi^{p}}_{T,{\rm!bc}}((\underline{\bm{u}}_{T},\underline{\rho}_{T},\underline{p}_{T});\cdot):\underline{\bm{V}}_{T}^{k}\to\mathbb{R}$ of the steady momentum equation, $sd^{\rm sms}_{T,{\rm!bc}}(\underline{\bm{u}}_{T},\underline{\rho}_{T};\cdot):\underline{Z}^{k}_{T}\to\mathbb{R}$ of the steady mass conservation equation and $sd^{\rm vol}_{T,{\rm!bc}}(\underline{\bm{u}}_{T};\cdot):\underline{Q}^{k+1}_{T}\to\mathbb{R}$ of the volume conservation equation are such that, for all $\underline{\bm{v}}_{T}\in\underline{\bm{V}}_{T}^{k}$, all $\underline{z}_{T}\in\underline{Z}_{T}^{k}$ and all $\underline{q}_{T}\in\underline{Q}_{T}^{k+1}$,
$sd^{{\rm slm},\pi^{p}}_{T,{\rm!bc}}\bigl((\underline{\bm{u}}_{T},\underline{\rho}_{T},\underline{p}_{T});\underline{\bm{v}}_{T}\bigr)=$ $-\frac{1}{2}\int_{T}\rho_{T}(\bm{\pi}_{T}^{p}\bm{u}_{T}\otimes\bm{u}_{T}):\nabla\bm{\pi}_{T}^{k}\bm{v}_{T}+\frac{1}{2}\int_{T}\rho_{T}(\bm{\pi}_{T}^{k}\bm{v}_{T}\otimes\bm{u}_{T}):\nabla\bm{\pi}_{T}^{p}\bm{u}_{T}$ (2)
$+\frac{1}{2}\int_{\partial T}\rho_{\partial T}\bigl(\bm{u}_{T}\cdot\bm{n}_{\partial T}\bigr)\left(\bm{\pi}_{\partial T}^{p}\bm{u}_{\partial T}\cdot\bm{\pi}_{T}^{k}\bm{v}_{T}-\bm{\pi}_{\partial T}^{k}\bm{v}_{\partial T}\cdot\bm{\pi}_{T}^{p}\bm{u}_{T}\right)$ (3)
$+\frac{1}{2}\int_{\partial T}\rho_{\partial T}\bigl|\bm{u}_{T}\cdot\bm{n}_{\partial T}\bigr|\left(\bm{\pi}_{T}^{p}\bm{u}_{T}-\bm{\pi}_{\partial T}^{p}\bm{u}_{\partial T}\right)\cdot\left(\bm{\pi}_{T}^{k}\bm{v}_{T}-\bm{\pi}_{\partial T}^{k}\bm{v}_{\partial T}\right)$ (4)
$+\int_{T}\mu\,{{\mathcal{G}}_{T}^{k}}\underline{\bm{u}}_{T}:{{\mathcal{G}}_{T}^{k}}\underline{\bm{v}}_{T}+\int_{\partial T}\frac{\mu}{h_{F}}\,\bm{\pi}_{\partial T}^{k}(\bm{u}_{T}-\bm{u}_{\partial T})\cdot\bm{\pi}_{\partial T}^{k}(\bm{v}_{T}-\bm{v}_{\partial T})$ (5)
@@
$sd^{\rm sms}_{T,{\rm!bc}}\bigl((\underline{\bm{u}}_{T},\underline{\rho}_{T});\underline{z}_{T}\bigr)=$ $-\frac{1}{2}\int_{T}\rho_{T}\bm{u}_{T}\cdot\nabla z_{T}+\frac{1}{2}\int_{T}z_{T}\bm{u}_{T}\cdot\nabla\rho_{T}$ (7)
$+\frac{1}{2}\int_{\partial T}\bigl(\bm{u}_{T}\cdot\bm{n}_{TF}\bigr)(\rho_{\partial T}{z}_{T}-{\rho}_{T}{z}_{\partial T})$ (8)
$+\frac{1}{2}\int_{\partial T}\bigl|\bm{u}_{T}\cdot\bm{n}_{TF}\bigr|({\rho}_{T}-\rho_{\partial T})({z}_{T}-{z}_{\partial T})$ (9)
@@
where $\bm{\pi}_{\partial T}^{\ell}$ is the $L^{2}$-orthogonal projector onto the broken polynomial space $\mathcal{P}^{\ell}(\partial T)^{d}$ defined as follows
$\mathcal{P}^{\ell}(\partial T)^{d}\coloneqq\left\{\bm{v}\in L^{2}(\partial T)^{d}\;:\;\text{$\bm{v}|_{F}\in\mathcal{P}^{\ell}(F)^{d}$ for all $F\in\mathcal{F}_{T}$}\right\},$
and $\bm{\pi}_{T}^{\ell}$ is the $L^{2}$-orthogonal projector onto $\mathcal{P}^{\ell}(T)^{d}$.
@@
Both in the momentum and the mass conservation equation, advection terms are treated with skew-symmetric formulations see (2)-(3) and (7)-(8), respectively, plus stabilization bilinear forms, see (4) and (9), respectively. We remark that, in the convective term discretization of (2)-(3)-(4), we consider $L^{2}$-projections of degree $k$ for test functions and $L^{2}$-projections of degree $p$, with p=kp{=}k or p=k+1p{=}k{+}1, for the advected velocity unknowns, leading to two different variants of the method. Note that the advective velocity is not projected. In the case p=k+1p{=}k{+}1, since $\bm{u}_{T}\in\mathcal{P}^{k+1}(T)$, it holds $\bm{\pi}_{T}^{p}\bm{u}_{T}=\bm{u}_{T}$, furthermore, since $\bm{u}_{\partial T}\in\mathcal{P}^{k}(\partial T^{{\rm i},{\rm D}})\subset\mathcal{P}^{k+1}(\partial T^{{\rm i},{\rm D}})$ and $\bm{u}_{\partial T}\in\mathcal{P}^{k+1}(\partial T^{{\rm N}})$, it holds $\bm{\pi}_{\partial T}^{p}\bm{u}_{\partial T}=\bm{u}_{\partial T}$, thus, contrary to the case p=kp{=}k, the formulations is not skew-symmetric. The choice p=kp{=}k, proposed in [4, Remark 3.7], leads to a provably stable formulation, see Sec. 3.7, while the choice p=k+1p{=}k{+}1, proposed in [8], provides improved accuracy in the convection-dominated regime. The diffusion term stabilization of (5), also involving $L^{2}$-projection operators, was first proposed by [20], see also [11]. Note that the HHO discretization of the mass conservation equation in (7) only relies on the cell velocity $\bm{u}_{T}$ which, as result of the HHO discretization of the volume conservation in (10), is $\bm{H}(\operatorname{div})$-conforming and divergence-free, see Sec. 3.5 for details.
@@
Let’s now take into account Dirichlet and Neumann boundary conditions. Given $(\underline{\bm{u}}_{T},\underline{\rho}_{T},\underline{p}_{T})\in\underline{\bm{V}}_{T}^{k}\times\underline{Z}_{T}^{k}\times\underline{Q}_{T}^{k+1}$, the (local) spatial discretizations $sd^{{\rm slm},\pi^{p}}_{T}((\underline{\bm{u}}_{T},\underline{\rho}_{T},\underline{p}_{T});\cdot):\underline{\bm{V}}_{T}^{k}\to\mathbb{R}$ of the steady momentum equation, $sd^{\rm sms}_{T}(\underline{\bm{u}}_{T},\underline{\rho}_{T};\cdot):\underline{Z}^{k}_{T}\to\mathbb{R}$ of the steady mass conservation equation and $sd^{\rm vol}_{T}(\underline{\bm{u}}_{T};\cdot):\underline{Q}^{k+1}_{T}\to\mathbb{R}$ of the volume conservation equation are such that, for all $\underline{\bm{v}}_{T}\in\underline{\bm{V}}_{T}^{k}$, all $\underline{z}_{T}\in\underline{Z}_{T}^{k}$ and all $\underline{q}_{T}\in\underline{Q}_{T}^{k+1}$,
$sd^{{\rm slm},\pi^{p}}_{T}\bigl((\underline{\bm{u}}_{T},\underline{\rho}_{T},\underline{p}_{T});\underline{\bm{v}}_{T}\bigr)$ $\coloneqq sd^{{\rm slm},\pi^{p}}_{T,{\rm!bc}}((\underline{\bm{u}}_{T},\underline{p}_{T});\underline{\bm{v}}_{T})-\frac{1}{2}\int_{\partial T^{{\rm D},{\rm N}}}\bigl(\bm{u}_{T}{\cdot}\bm{n}_{\partial T}\bigr)\bigl(\rho_{\partial T}\bm{\pi}_{\partial T}^{p}\bm{u}_{\partial T}{\cdot}{\bm{\pi}_{\partial T}^{k}}\bm{v}_{\partial T}\bigr)$
@@
$+\int_{\partial T^{\rm N}}\left(\bigl(\bm{u}_{\partial T}{\cdot}\bm{n}_{\partial T}\bigr)^{\oplus}\rho_{\partial T}+\bigl(\bm{u}_{\partial T}{\cdot}\bm{n}_{\partial T}\bigr)^{\ominus}g_{\rm In}\right)\bigl({\bm{\pi}_{\partial T}^{p}}\bm{u}_{\partial T}{\cdot}{\bm{\pi}_{\partial T}^{k}}\bm{v}_{\partial T}\bigr)$
@@
$+\int_{\partial T^{\rm D}}\left(\bigl(\bm{g}_{\rm D}\cdot\bm{n}_{\partial T}\bigr)^{\oplus}\rho_{\partial T}+\bigl(\bm{g}_{\rm D}\cdot\bm{n}_{\partial T}\bigr)^{\ominus}g_{\rm In}\right)\;(\bm{g}_{\rm D}\cdot\bm{\pi}^{k}_{\partial T}\bm{v}_{\partial T})$